OpenStudy (anonymous):

Write the quadratic equation whose roots are −4 and −6 , and whose leading coefficient is 4 .

OpenStudy (anonymous):

what

OpenStudy (cj49):

x = -4 and x = -6 so x + 4 = 0 and x + 6 = 0 (x + 4)(x + 3) = 0 x^2 + 3x + 4x + 12 = 0 x^2 + x + 12 = 0 Multiplying by 4 4x^2 - 4x - 48 = 0

OpenStudy (cj49):

4x^2 + 4x + 48 = 0

OpenStudy (xapproachesinfinity):

well since the zeros are -4 and -6 then you quadratic function is written like this \(\Large \color{midnightblue}{f(x)=a(x+6)(x+4)}\) a is given to you to be 4 so \(\Large \color{midnightblue}{f(x)=4(x+6)(x+4)}\) you question didn't really as if we should write it in standard form so i suppose you just leave that way

OpenStudy (xapproachesinfinity):

ask*